A 0.5m3 rigid tank contains refrigerant at 180C. If saturated liquid occupies 10% of the volume, determine (a.) the total mass of refrigerant, (b.) the quality of the saturated mixture.
Answer #1

At 180°C

vf = 0.001190 m^3/kg and

vg = 0.08609 m^3/kg

saturated liquid occupies 1/10 of volume and rest vapour form 9/10 of volume

a)

mf = Vf / vf = [(1/10)*(0.5 m^3)] / 0.001190 m^3/kg

mf = 42 kg

mg = Vg / vg

mg = [(9/10)*(0.5 m^3)] / 0.08609 m^3/kg

mg = 5.23 kg

mt = mf + mg

mt = 42 kg + 5.23 kg

mt = 47.23 kg

b)

x = mg/mf

x = 5.23 kg / 42 kg

x = 0.124523